Suppose it is known that among all public colleges and universities, 23.8% of them are  planning to move their summer 2020 classes online. If a simple random sample of 182 public colleges and universities is selected and the proportion   of this sample who are planning to move their summer 2020 classes online are determined, if appropriate describe completely the sampling distribution of
Answer #1

Solution

Given that,

p = 0.238

1 - p = 1 - 0.238 = 0.762

n = 182

= p = 0.238

  [p ( 1 - p ) / n] = [(0.238 * 0.762) / 182 ] = 0.0316